OpenStudy (anonymous):

Medal for best answer!! Martin combined different amounts of orange juice, pineapple juice, and strawberry juice to make 1850 milliliters (ml) of fruit punch. The amount of orange juice combined was x ml. The amount of pineapple juice combined was 100 ml more than the amount of orange juice, and the amount of strawberry juice was half the amount of orange juice. This is represented in the equation below. x + (x + 100) + one half x = 1850 What was the difference in the amounts of pineapple juice and strawberry juice Martin combined? 100 ml 450 ml 700 ml 950 ml

OpenStudy (calculusxy):

Will answering this question help u if I help you solve for the value of x?

OpenStudy (anonymous):

Yes

OpenStudy (calculusxy):

Okay then.

OpenStudy (calculusxy):

When combining the different amounts of x, we usually count them and then say 'x' next to it. For example, if there was x+x+x+x, we would write that as 4x. Using this example, can you tell me what should be written for the number of x's (eg. 4x).

OpenStudy (anonymous):

First I am not stupid I know about combining like terms and about x + x + x + x = 4x but we can't really combine these x's because they are for different things like the first x is for orange juice the second x is for orange juice but you are adding 100 to it for the pineapple juice and the last x is for the strawberry juice.

OpenStudy (calculusxy):

I really didn't mean that you are not smart. I am sorry if you felt offended. But since we are trying to find the value of x (orange juice), you can combine like terms. Don't worry about what the variables mean.

OpenStudy (anonymous):

1 1/2x + (x + 100) = 1850 The only reason I am not combining the x in ( ) is because it has to get added to 100 so shouldn't we leave that alone but if not then here 2 1/5x + 100 = 1850

OpenStudy (anonymous):

And what you said is ok I just miscommunicated what you had said, it is ok though.

OpenStudy (calculusxy):

\[x+(x+100)+0.5x=1850\] take out the parenthesis and then combine the like terms. \[2x+100+05x\] combine 2x+0.5x to get \[2.5x+100=1850\] now solve for x.

OpenStudy (anonymous):

The equation that you wrote is the same as the last one I wrote I just wasn't sure. x = 700

OpenStudy (calculusxy):

Okay but that's not done yet

OpenStudy (anonymous):

My answer is 450 because after I substitute everything in my pineapple juice equaled 800 and my strawberry juice equaled 350. subtract those and you get 450

OpenStudy (calculusxy):

You are correct. :)

OpenStudy (anonymous):

Can you help me with one more?

OpenStudy (calculusxy):

Sure I will try my best :)

OpenStudy (anonymous):

What should be measured to calculate the average rate at which water flows in a dam? Distance of water flow and duration of water flow Number of liters of water flowing and duration of water flow Distance of water flow and size of the dam from which water flows Number of liters of water flowing and size of the dam from which water flows

OpenStudy (calculusxy):

I believe that it is the distance of water flow and duration of water flow,

OpenStudy (anonymous):

I was thinking that at first but I wasn't sure. Thanks :)

OpenStudy (calculusxy):

No problem. If you like my answer, you can be my fan or give me a medal :)

OpenStudy (anonymous):

Ok lol